If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?

Answer by scott8148(6628)

they each do a FRACTION of the job, based on their individual times ___ the fractions add to 1 (the WHOLE job)

let x=combined time

x/4 + x/6 + x/2 = 1

multiplying by 12 (LCD) ___ 3x + 2x + 6x = 12

11x = 12 ___ x = 12/11